A population of values has a nornal distribution with μ sample of size n = 41. 214 and σ 22.5. You intend to draw a random Find the probability that a single randomly selected value is between 208.4 and 223.1. P(208.4<X< 223.1)- Find the probability that a sample of size n 41 is randomly selected with a mean between 208.4 and 223.1. P(208.4< M< 223.1)- Enter your answers as numbers accurate to 4 decimal places. Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted Points possible: 1 License Unlimited attempts. Submit
In a recent year, the Better Business Bureau settled 75% of complaints they received. (Source: USA Today. March 2, 2009) You have been hired by the Bureau to investigate complaints this year involving computer sto ou plan to select a random sample of complaints to estimate the proportion of complaints the Bureau is able to settle. Assume the population proportion of complaints settled for the computer stores is the 0.75, as mentioned above. Suppose your sample size is 177. What is the probability that the sample proportion will be within 9 percent of the population proportion? Note: You should carefully round any z-values you calculate to 4 decimal places to match wamaps approach and calculations. Answer - Points possible: 1 Enter your answer as a number accurate to 4 decimal places.) License Unlimited attempts.
Answer #1

1)a)

for normal distribution z score =(X-μ)/σx
a)
b)here mean=       μ= 214
std deviation   =σ= 22.5000
probability = P(208.4<X<223.1) = P(-0.25<Z<0.4)= 0.6571-0.4017= 0.2553

b)

sample size   =n=

41
std error=σ=σ/√n= 3.5139
probability = P(208.4<X<223.1) = P(-1.59<Z<2.59)= 0.9952-0.0555= 0.9397

2)

for normal distribution z score =(p̂-p)/σp
here population proportion=     p= 0.750
sample size       =n= 177
std error of proportion=σp=√(p*(1-p)/n)= 0.0325
probability = P(0.66<X<0.84) = P(-2.77<Z<2.77)= 0.9972-0.0028= 0.9944